Identify of an alkane gas that has a density of 1.16 g/l measured at pressure of 0.984 atm and temperature of 37C?
1
Expert's answer
2012-12-03T08:58:25-0500
PV = nRT
n = PV/RT
n = m/Mw
m/Mw = PV/RT
Mw = mRT/PV = 1.16*8.31*(273+37)/(1*(101.3*0.984) = 30

CxHy

30/12=2..... x=2
30-2*12=6 y=6

Answer: C2H6